The area of a rectangle is 156 feet. What are the demensions of the rectangle if the length is one foot longer than the width?

Answer by jim_thompson5910(35256)

The area of a rectangle is 156 feet. What are the demensions of the rectangle if the length is one foot longer than the width?


Let x be the width. Since the "length is one foot longer than the width", we know that the length is x+1

Area = Length * Width

A = (x+1)*x ... plug in x+1 for the length, x for the width

A = x(x+1)

156 = x(x+1) ... plug in the given area

156 = x^2 + x

0 = x^2 + x - 156

x^2 + x - 156 = 0

(x + 13)(x - 12) = 0 ... note: you can use the quadratic formula if you don't see how it factors

x + 13 = 0 or x - 12 = 0

x = -13 or x = 12

Since we cannot have a negative width, toss out x = -13. So the only solution is x = 12

The width is 12 feet

The length is x + 1 = 12+1 = 13 feet